OpenStudy (anonymous):

Find the inverse of the function. If the function does not have an inverse, write NO INVERSE. f(x) = x – 3

OpenStudy (anonymous):

y = x- 3 x = y - 3 x +3 = y y = x + 3

OpenStudy (anonymous):

F(x)=y y=x-3 Switch x and y x=y-3 Solve for y -y=-x-3 y=x+3

OpenStudy (anonymous):

thank you